Problem 11-2A (Part Level Submission) The stockholders’ equity accounts of Whispering Winds Corp. on January 1, 2017, were as follows. Preferred Stock (7%, $100 par noncumulative, 5,000 shares authorized) $300,000 Common Stock ($4 stated value, 300,000 shares authorized) 1,000,000 Paid-in Capital in Excess of Par Value—Preferred Stock 15,000 Paid-in Capital in Excess of Stated Value—Common Stock 480,000 Retained Earnings 699,500 Treasury Stock (5,000 common shares) 40,000 During 2017, the corporation had the following transactions and events pertaining to its stockholders’ equity. Feb. 1 Issued 5,000 shares of common stock for $30,000. Mar. 20 Purchased 1,000 additional shares of common treasury stock at $8 per share. Oct. 1 Declared a 7% cash dividend on preferred stock, payable November 1. Nov. 1 Paid the dividend declared on October 1. Dec. 1 Declared a $0.70 per share cash dividend to common stockholders of record on December 15, payable December 31, 2017. Dec. 31 Determined that net income for the year was $277,400. Paid the dividend declared on December 1. (b) Enter the beginning balances in the accounts and post the journal entries to the stockholders’ equity accounts. (Post entries in the order of journal entries posted in the previous part. For accounts that have zero ending balance, the entry should be the balance date and zero for the amount.)
Answer #1

Preferred stock

Jan 1 Balance $300000
Dec 31 Balance $300000

Paid in capital in excess of Par value- Preferred stock

Jan 1 Balance $15000
Dec 31 Balance $15000

Common Stock

Jan 1 Balance $1000000
Feb 1 20000
Dec 31 Balance $1020000

Paid in capital in excess of Stated value- Common stock

Jan 1 Balance $480000
Feb 1 10000
Dec 31 Balance $490000

Retained Earnings

Dec 31 (21000+174300)= $195300 Jan 1 Balance $699500
Dec 31 277400
Dec 31 Balance $781600

Treasury Stock

Jan 1 Balance $40000
Mar 20 8000
Dec 31 Balance $48000

Cash Dividends

Oct 1 (300000*7%)= $21000
Dec 1 (1000000/4-5000+5000-1000)*0.70= 174300 Dec 31 $195300
Dec 31 Balance 0